OpenStudy (anonymous):

What expression is the simplified form of the square root of y to the fourth power? y y2 y times the square root of y y squared times the square root of y

OpenStudy (michele_laino):

\[(\sqrt{y})^{4}=y ^{4/2}=y^{2}\]

OpenStudy (michele_laino):

because: \[\sqrt{y}=y ^{1/2}\] so the second option

OpenStudy (anonymous):

Thank you!

OpenStudy (michele_laino):

Thank you!

OpenStudy (anonymous):

Can you check to see if I got something right?

OpenStudy (anonymous):

Which expression is the simplified form of the square root of 49 times a to the fourth power? 7a 7 times a times the square root of a 7a2 7 times a squared times the square root of a

OpenStudy (anonymous):

I got D

OpenStudy (anonymous):

that is incorrect

OpenStudy (anonymous):

you should use latex to write your answers because its unclear what the question is and what the answers are

OpenStudy (anonymous):

or at least use parenthesis

OpenStudy (anonymous):

$$\sqrt{49a^{4}}$$

OpenStudy (anonymous):

is that your question?